# Sea-Land-Cloud Segmentation in Satellite Hyperspectral Imagery by Deep Learning

### v2

@@ 4.1.3 Segmentation Performance (III): False Detection Ratios, Data Loss, and Downlink Inefficiency
Considering the initial analysis focused solely on data loss and the subsequent analysis regarding the trade-off between both data loss and downlink efficiency, we next provide our conclusion. On the one hand, the 2D-Justo-UNet-Simple, with approximately 7 thousand parameters, minimizes the loss of valuable data for most classes. The only exception is the sea class, where it incurs an acceptable 5% data loss in contrast to the 1% loss by 2D-UNet-FAUBAI-reduced which uses around 2 million parameters and hence the FAUBAI’s network is less suitable for on-board inference. On the other hand, when we equally consider both data loss and downlink usage efficiency, our model, 1D-Justo-LiuNet, excels in minimizing data loss, achieving low rates of 2%, 5%, and 3%, while simultaneously demonstrating relatively minimal inefficiency rates of 2%, 3%, and 15% for sea, land, and cloud classes, respectively. The findings suggest that when employing 1D-Justo-LiuNet, downlink efficiency slightly deteriorates due to an under-detection of clouds, leading to an overabundance of 15% of cloud data that would be downlinked from the satellite. We attribute the finding to the dataset’s class imbalance, where the cloud class constitutes the minority. Nevertheless, the over-detection of the majority classes sea (2%) and land (3%) pixels is smaller, thereby increasing the overall efficiency of the downlink channel.
4.1.4 Segmentation Performance (IV): Priority Rankings for Downlink
As part of the further evaluation of the models’ performance, we conduct an additional analysis of the ranking system to evaluate its effectiveness in sorting HS images, as explained next. We center our analysis on the Spearman’s coefficient (Table II) to determine which model precisely ranks the images based on the sea, land, and cloud coverage levels calculated from the segmented images. According to Tables II and LABEL:Table:models_sorted_according_to_performance, the experimental results confirm that 1D-Justo-LiuNet and 1D-Justo-HuNet are the top performers, surpassing again the state-of-the-art models. As a matter of fact, Table II shows that 1D-Justo-LiuNet and 1D-Justo-HuNet are the only models capable of reaching the ideal Spearman’s coefficient of 1.00, indicating the perfect sorting of the images in the rankings, with no errors compared to the ground-truth ranks for sea, land, and cloud levels. This emphasizes the selection of 1D-Justo-LiuNet (or possibly 1D-Justo-HuNet) as the preferred choice for the proposed on-board ranking system.
Furthermore, the results reveal that our 2D model, 2D-Justo-UNet-Simple, although not achieving the same level of ranking performance as the 1D models, also surpass most of their fellow 2D models from the state of the art. Nonetheless, our model, 1D-Justo-LucasCNN, does not produce improved results, as it only yields the same figures to those achieved by other state-of-the-art models (1D-ML/SGD, 2D-CUNet Reduced, and 2D-UNet FAUBAI).
4.1.5 Segmentation Performance (V): Model Generalisation in Deployment Set
Aligning with the segmentation literature for satellite imagery [13], we conduct generalization tests in the unseen on-ground deployment set of 30 images, relying on expert visual inspection (similar to [13]) as this set has no ground-truth labels available. From the generalization tests, we next illustrate the segmentation results for one of the deployment images, specifically a HYPSO-1 capture taken over the Namib desert in Africa [34], as depicted in Fig. 6. Fig. 6(a) presents the segmentation results obtained using 1D-Justo-LiuNet. The figure includes a grayscale L1b radiance image at a wavelength of $\lambda$=699.61 nm, followed by the model’s segmented image, a confidence probability map indicating the model’s certainty for each pixel-level prediction, and summarised information regarding sea, land, and cloud coverage levels used to sort the images in the ranking system. The figure illustrates that 1D-Justo-LiuNet performs adequately, with a slightly reduced certainty in class transitions while maintaining a high level of overall certainty in its predictions. Regarding 2D-CNN models, due to the limited generalization capability observed in the deployment set for 2D-Justo-UNet-Simple (which is, however, a top-performer in the test set), we choose to conduct generalization tests on other 2D-CNN models of interest instead, such as 2D-CUNet++ Reduced, as illustrated in Fig. 6(b). The figure highlights that 2D-CUNet++ Reduced encounters important misclassifications at the non-overlapping patch borders, consistent with the confidence observed in the patch edges in addition to the decreased certainty in the class transitions. This phenomenon may be attributed to a variety of factors, including the network’s padding of input patches, among other potential causes.
In the supplementary material for the generalisation tests on the rest of the deployment set, we provide the outcomes for the remaining deployment images. From the overall on-ground deployment, we conclude that 1D-Justo-LiuNet generalizes acceptably for our needs. However, its 1D-CNN architecture can result in predictions with higher entropy, leading to less smooth class transitions. This is attributed to the absence of spatial context in its pixel-level predictions. On the other hand, 2D-CUNet++ Reduced, while still generalizing effectively, exhibits higher misclassifications at patch borders. Nevertheless, it remarkably provides predictions with lower entropy, resulting in smoother class transitions, due to the utilization of spatial neighboring information during pixel-level predictions.
4.1.6 Segmentation Performance (VI): Summary
In light of the previous sections, in addition to the key finding that 1D models outperform their 2D counterparts in our experiments, we conclude that the 1D-Justo-LiuNet model consistently outperforms the current state-of-the-art models, all trained for the same sea-land-cloud segmentation task in HYPSO-1 L1b calibrated data. Specifically, the 1D-Justo-LiuNet surpasses the literature in terms of segmentation performance, assessed by average accuracy and false detection ratios trade-off. Furthermore, our results for the 1D-Justo-LiuNet not only surpass the performance reported by the soil spectroscopy literature [45], but our findings also exhibit remarkable consistency, even within the context of the ranking system application case (Spearman’s coefficient), where the 1D-Justo-LiuNet also outperforms the other models. Additionally, 1D-Justo-LiuNet demonstrates acceptable performance when subjected to generalization tests on the unseen on-ground deployment set. However, we stress that our results and conclusions are specific to the sea-land-cloud dataset [34], and we acknowledge that model performance may differ when the data source is altered in any manner. For instance, a model that performs well on one dataset may not excel on another. To facilitate testing of the models used in this work on other data sources, we provide the software code for all models in the supplementary materials accompanying this paper.
4.1.7 Parameter Count
Following the analysis of the models’ performance, we next examine the number of model parameters (Table II) that need to be uplinked to a satellite for in-orbit deployment. Our lightweight model 1D-Justo-LiuNet surpasses the existing literature in terms of parameter count ranking at the top (Table LABEL:Table:models_sorted_according_to_performance), as the model with the fewest parameters (4,563), even when compared to other lightweight models in the literature, such as 2D-CUNet, 2D-CUNet++, 2D-CUNet Reduced, and 2D-CUNet++ Reduced. Hence, the lightweight 1D-Justo-LiuNet model stands out as the most practical choice when solely considering the reduction in on-board memory required for storing the model parameters. Moreover, despite some of our other models, such as 1D-Justo-HuNet and 2D-Justo-UNet-Simple, exhibit a slight increase in the number of parameters compared to 1D-Justo-LiuNet, they still maintain their parameter count below ten thousands, which is still lower than the number of parameters of the other state-of-the-art models. As mentioned earlier, the lowest-ranked 2D-UNet FAUBAI network utilizes over 26 million parameters. Even its reduced version, 2D-UNet FAUBAI reduced, has nearly 2 million parameters, which we also find excessive for on-board deployment while it does not improve the model’s performance neither in terms of segmentation accuracy nor in terms of image ranking. Finally, we emphasize that a significant majority of our top models in terms of performance (accuracy and Spearman’s coefficient in Table LABEL:Table:models_sorted_according_to_performance), also occupy the best ranks in terms of parameter count. Therefore, we can assert that we achieve our objective outlined in Section 2 regarding surpassing the state-of-the-art performance, while significantly compressing the models.
4.1.8 Inference Time
The final metric of our analysis is the average inference time per HS image (Table II). Despite the strengths of the 1D models, with particular emphasis on the favorable results by the 1D-Justo-LiuNet, 1D models exhibit a drawback of extended inference times on the processing architecture used in the tests, even considering their low parameter count. In Table II, we observe that both 1D ML and DL models, especially DL models with more intrinsic complexity, exhibit lengthier inference times. In fact, the results reveal that the 1D-Justo-LiuNet has the lengthiest inference time, at around 15 seconds for a single HS image. The results demonstrate, however, that 2D models significantly decrease the inference time, reducing it to the order of milliseconds. The disparity in speed may be attributed due to the potentially greater number of floating-point operations per second (FLOPS) required by the 1D models, among other factors affecting the inference time.
Although satellites like HYPSO-1 experience data latency lasting minutes to hours [29], we propose the following solution if inference times in the order of seconds (e.g., by 1D-Justo-LiuNet) are perceived as lengthy for in-orbit inference. Employing the HYPSO-1 satellite as an example case, its On-board Processing Unit (OPU) integrates an FPGA System-on-Chip (Zynq-7030/XC7Z030-1) [10]. The system features an ARM Cortex-A9 processor and provides superior FPGA logic resources in comparison to other platforms, such as the Zynq-7020, previously employed in cloud detection within satellite imagery [14]. As future work, we suggest implementing the 1D-Justo-LiuNet in the processor side to confirm whether the inference time is still lengthy. We anticipate that the processor’s lower-level software implementation may yield shorter inference times compared to running inference in higher-level languages like Python. If the inference time would still remain too long, we recommend a combined software-hardware co-design approach. This approach accelerates computationally intensive tasks for faster inference, while maintaining low power consumption, by using the FPGA fabric logic.
4.2 Results Analysis for Models Trained on 3 Channels
@@
We finalize the discussion section by addressing the impact of reducing the number of channels from 112 to 3 on the reported metrics. As indicated in Table IV, when we train the models using the 3-channel dataset instead, the inference times decrease compared to the models trained with 112 channels, as there is now less data to process. Moreover, there is a reduction in the number of model parameters due to the smaller number of components within each convolutional kernel when computing the respective 2D convolutions across the reduced spectral dimension. Nevertheless, the reduction in feature channels leads to a decreased segmentation performance, with the highest accuracy reaching 0.87, unlike the 0.93 top accuracy for the models trained with 112 channels. The decline in performance is also evident in the evaluation of the ranking system, for which none of the models achieve the ideal coefficient of 1.00, as observed earlier with 112 channels. As a matter of fact, the highest Spearman’s coefficient for 3 channels is notably lower at 0.83.
5 Conclusion
In this work, we conducted experiments including the training and comparison of 16 machine learning and deep learning models, whose software codes are provided within this article along with additional supplementary materials, for future in-orbit deployment using the Sea-Land-Cloud-Labeled dataset from the HYPSO-1 mission [34], as an illustrative case. Our experiments demonstrated that when segmenting sea, land, and cloud categories in hyperspectral satellite imagery, 1D models surpass 2D models in terms of performance. This conclusion also holds for our proposed ranking system combining priority downlink according to sea, land and cloud coverage. Our proposed lightweight models, such as 1D-Justo-LiuNet and 1D-Justo-HuNet, consistently surpass the performance of state-of-the-art models. Moreover, they demonstrate remarkable model compression, reaching up to 90.50% when compared to the existing models. This makes 1D-Justo-LiuNet and 1D-Justo-HuNet well-suited for future in-flight deployment. Furthermore, we also found that using L1b calibrated radiance data, instead of raw sensor data, improved segmentation performance. Moreover, training models with only three spectral channels led to shorter inference times and reduced parameter counts, at the cost of lower segmentation performance. Finally, as future work, we suggest to study the implementation of processing architectures that are used aboard satellites and conducting experimental tests to deploy, for instance, the 1D-Justo-LiuNet model in orbit.
Declaration of Competing Interest
The authors declare that they have no known competing interests that could have appeared to influence the work reported in this paper.
@@ 5 Conclusion
To begin, the labeled dataset used in this paper is openly available to download at https://ntnu-smallsat-lab.github.io/hypso1_sea_land_clouds_dataset/. Our supplementary material is openly available at https://github.com/NTNU-SmallSat-Lab/s_l_c_segm_hyp_img/. We provide the following resources. Firstly, the Python model codes accompanied by a notebook that illustrates how to perform inference on HYPSO-1 data, by 1D-CNN and 2D-CNN models. Secondly, there is an executable software program designed to serve as an oceanic-terrestrial-cloud segmenter for hyperspectral imagery acquired by the HYPSO-1 satellite. Thirdly, we provide trained models with learned parameters, with 46 distinct configurations, for the HYPSO-1 data. Fourth, we offer additional in-depth architectural details for the models. Fifth, we include information regarding the training, validation, and test data splits. Sixth, we share an extensive set of experimental figures, including multiple metrics. Seventh, we provide a deployment set consisting of 30 unlabeled HYPSO-1 images. Lastly, we present the generalisation tests in the complete deployment set.
Acknowledgements
The research leading to these results has received funding from the NO Grants 2014 – 2021, under Project ELO-Hyp contract no. 24/2020, and the Research Council of Norway grant no. 328724 (Green Platform) and no. 325961 (HYPSCI). We extend our gratitude to Anum Masood, Per Gunnar Kjeldsberg, and Simen Netteland.
@@ Appendix A Direct Channel Reduction
Considering that the legacy algorithms for cloud detection in HS satellite imagery typically select 3 channels to mimic RGB image processing [11, 12, 13, 14], we also obtain another dataset by reducing the 112 channels to merely 3 channels. To achieve this reduction, we explore the literature of dimensionality reduction extensively studied for HS satellite missions [46, 47, 48]. First, the focus on reducing dimensions originates from the importance of having low-dimensional data which can save computation resources, memory usage and power. Next, some of the techniques proposed by the literature include methods like Principal Component Analysis (PCA) [49] and Independent Component Analysis (ICA) [50]. In alignment with the state-of-the-art techniques, we employ PCA to reduce the 112 channels (normalized) to 3. To alleviate to a certain extent the computational demands of satellites operating often with tight power constraints, we propose to conduct both the training and inference of PCA exclusively on the ground. Eliminating PCA inference on the satellite enables potential power savings, which could be of importance to conduct the subsequent segmentation inference during flight. Consequently, and as also outlined by the literature [12], we suggest implementing on board a direct channel selection (3 channels) instead of deploying PCA inference. In this work, the purpose of employing PCA only on ground is to identify in advance the channels containing the most useful information in the data. In particular, we use PCA to identify the channels that have the greatest influence on a single principal component, as determined by the weight loadings of those channels. In the case of the 112-channel L1b dataset, we find that the majority of influential channels are predominantly within the NIR range. However, given the necessity of considering other spectral ranges in addition to NIR for the adequate segmentation of sea, land, and clouds – as the NIR range alone is likely to be insufficient – we opt to pick the top three influential channels from only the visible blue, second the visible green and red combined, and finally the NIR. In the dataset, the most influential channels within each spectral range are: 412.72 nm (channel 7) for the blue spectrum, 699.61 nm (channel 89) in the green and red spectra, and 747.77 nm (channel 103) in the NIR spectrum. These align closely with the L1b radiance bands frequently analyzed during HYPSO-1 operations, especially for channels 89 and 103.

### v3

@@ IV-A5 Class Segmentation Performance: False Detection Ratios, Data Loss, and Downlink Inefficiency
In summary, when evaluating segmentation performance for each individual class considering a reasonable trade-off between data loss and downlink efficiency, 1D-Justo-LiuNet gnerally outperforms the other state-of-the-art models. We remark that 1D-Justo-LiuNet also excels in minimizing data loss, achieving acceptable low rates of 2%, 5%, and 3%, while simultaneously demonstrating relatively minimal inefficiency rates of 2%, 3%, and 15% for sea, land, and cloud classes, respectively. On the one hand, these findings suggest that downlink efficiency slightly deteriorates due to an under-detection of clouds, leading to an overabundance of 15% cloud data that would be downlinked from the satellite. Despite not critical, we attribute this to the dataset’s class imbalance, where the cloud class constitutes the minority. On the other hand, the over-detection of the majority classes sea (2%) and land (3%) pixels is smaller, thereby increasing the overall efficiency of the downlink channel.
Before concluding the discussion on model performance, we compare our results with other state-of-the-art methods for HS segmentation. For example, a Support Vector Machines (SVM) model was previously tested and deployed in orbit for HS segmentation on HYPSO-1 [18]. The SVM achieved a maximum average accuracy of 0.85, with a significantly lower minimum accuracy of 0.19. In contrast, 1D-Justo-LiuNet demonstrates a higher accuracy of 0.93, while the lowest accuracy, by 2D-CUNet, is 0.72, improving overall performance on HYPSO-1.
IV-A6 Generalization Performance in HYPSO-1 Deployment Set
Up to this point, we have reported results based on the test set, where metrics such as accuracy and Spearman’s coefficient could be calculated using ground-truth labels. However, to evaluate the model on a larger, unseen dataset, we utilize a deployment set, which lacks ground-truth. As a result, the further generalization assessment in HYPSO-1 is conducted through expert visual inspection, a common practice in the literature [13]. Given the promising results of lightweight 1D-CNNs for future edge inference, we focus on their generalization capabilities compared with 2D-CNNs, which have fewer parameters than ViTs, making them more suitable for on-board inference. When comparing the segmentation results of 1D-CNN models in Figs. 6-8 with the results from 2D-CNNs in Figs. 9-17, we confirm that 1D-CNNs outperform 2D-CNNs, with 1D-Justo-LiuNet achieving the best performance (Fig. 6).
Regarding the sea, land, and cloud coverage levels shown in the pie charts in the same figures, a cloud coverage level of around 30% is reasonable to expect for this scene. The 1D models predict this level accurately, while the 2D models often yield significantly different results, indicating much higher cloud coverage. This discrepancy can lead to the irrevocable loss (or degradation) of valuable surface data. Although 2D-CUNet (Fig. 11) and 2D-CUNet++ (Fig. 12) may provide acceptable cloud coverage levels, they misclassify some ground data as clouds, still resulting in irrevocable loss of information. Thus, 1D-CNNs such as 1D-Justo-LiuNet remain the most suitable option in terms of performance. Even when considering the model’s confidence in terms of probability shown in the figures, 1D models show more certainty in pixel-level predictions compared with 2D models. As a matter of fact, 1D-Justo-LiuNet (Fig. 6), clearly stands out with the highest confidence and almost no uncertainty, whereas the remaining models, especially the 2D ones, exhibit significantly lower certainty.
Furthermore, we note that for all the models, feature extraction is challenging at class boundaries, with lower certainty at these transitions. This issue may increase the risk of higher misclassification, for instance, in images with high-entropic class distributions or scenes with high variability. We attribute this to the fact that 1D models do not consider spatial neighboring context, while 2D models may still require additional training and regularization to improve performance at class transitions. Additionally, we observe that 1D models generally produce predictions with higher entropy. This is because they lack spatial context unlike 2D models, leading to relatively less smooth transitions between predicted classes. Finally, Figs. 9-17 show that 2D models tend to misclassify pixels at the edges of the non-overlapping patches, exhibiting reduced certainty in these tile regions. This issue may arise from several factors, including the padding effect of the network on the edges of each input patch being processed. Based on the overall analysis of generalization in HYPSO-1, we conclude that 1D-Justo-LiuNet remains the top-performing model for future in-orbit deployment.
IV-A7 Models Trained on 3 Channels
@@
We also address the impact of reducing the number of channels from 112 to 3 on the reported metrics. As indicated in Table IV, when we train the models using the 3-channel dataset from HYPSO-1, the inference times decrease compared with the models trained with 112 channels, as there is now less data to process. Moreover, there is a reduction in the number of model parameters due to the smaller number of components within each convolutional kernel when computing the respective 2D convolutions across the reduced spectral dimension. However, the reduction in feature channels leads to a decreased segmentation performance, with the highest accuracy reaching 0.87, unlike the 0.93 top accuracy for the models trained with 112 channels. The decline in performance is also evident in the evaluation of the ranking system, for which none of the models achieve the ideal Spearman’s coefficient of 1.00, as observed earlier with 112 channels. As a matter of fact, the highest coefficient for 3 channels is notably lower at 0.83. In conclusion, to achieve higher segmentation performance in HS imagery, inference should be conducted using most spectral channels due to the richness of information in this domain. Given the high dimensionality of the data, to reduce complexity, 1D-CNNs should be utilized, which also boost segmentation performance over 2D-CNNs. Although Table IV does not include results for 1D-CNNs and FastViTs using only 3 channels, we perform these experiments on the RGB SID dataset.
IV-A8 Knowledge Distillation
In an attempt to boost the performance of 1D-Justo-LiuNet, we conduct additional experiments using knowledge distillation, with FastViT-S12 as the teacher model and 1D-Justo-LiuNet as the student. However, we do not obtain any performance improvements, i.e., the accuracy of 1D-Justo-LiuNet remains at 93%. This can be explained by the fact that, as shown in Table II, 1D-Justo-LiuNet has a higher performance level (0.93) than the FastViT-S12 teacher (0.90).
IV-A9 The Effect of L1b Radiance Calibration
Before concluding the discussion of the main results, we address the impact of L1b radiance calibration on segmentation performance. Table II shows models trained on L1b calibrated radiance, a standard format in HSI. Radiance not only allows satellite comparisons but also removes sensor-specific characteristics and biases. We trained our models in radiance to ensure that the results can be generalized to satellites different from HYPSO-1. Beyond this, it is of interest to assess how L1b radiance calibration affects segmentation performance to determine whether in-orbit inference should occur before or after calibration.
For brevity, we provide additional comparative tables in the supplementary material showing segmentation results when models are trained for pre- and post-calibration inference. The highest-performing 1D-CNN models show slightly better performance with L1b radiance than with raw data. A similar trend is observed with top 2D-CNNs, which also perform marginally better with radiance. However, lower-performing models tend to perform better with raw data. While the top-performing models benefit slightly from radiance calibration, the improvement is minimal and may not justify the computational cost of performing calibration in orbit. In any case, we attribute the marginal improvement to calibration correcting errors and biases, allowing top-performing models to focus on inherent data patterns rather than overfitting on sensor-induced characteristics, thus improving generalization. However, we recommend testing both calibrated and raw sensor data when training models on different datasets to verify if calibrated data, such as radiance or reflectance, continues to enhance performance.
IV-B Discussion of Generalization Results
To validate the generalization of our results and conclusions, we use NASA’s EO-1 dataset from the Hyperion HS sensor. Our goal is to determine whether 1D-Justo-LiuNet continues to outperform 2D-CNNs, as observed in the HYPSO-1 dataset. As shown in Table V, we consider two 2D-CNNs. The first one is 2D-Justo-UNet-Simple, while the second one is nnU-Net 2D, a recent model that has shown promising results. The latter deep network [51] is a self-configuring nnU-Net, which was later compressed [52] and presented in the competition On Cloud N: Cloud Cover Detection Challenge. After training, validating, and testing the 1D-CNN and 2D-CNN models on the EO-1 dataset, Table V confirms that 1D-Justo-LiuNet once again outperforms the 2D-CNNs. However, the table shows that this conclusion applies only when most channels are used. When the number of channels is drastically reduced, the spectral information loss is significant, and as expected, 1D-Justo-LiuNet does not outperform 2D models. However, the results indicate that the best segmentation performance is only achieved by 1D-Justo-LiuNet, with an accuracy of 0.82 and an F1 score of 0.97, particularly when a larger set of spectral channels is utilized. Therefore, we conclude that optimal segmentation in HS imagery is achieved when using 1D-CNNs, such as 1D-Justo-LiuNet, in conjunction with the largest possible spectral range. In short, our conclusions on HYPSO-1 are confirmed. A detailed description of our experiments on the NASA’s EO-1 Hyperion dataset is provided in [37].
Finally, Table VI presents results from the RGB Satellite Imagery of Dubai (SID) dataset. The results on this 3-channel dataset indicate that our earlier conclusions generalize only to the hyperspectral domain. As expected, for satellite RGB imagery, 2D-CNNs and Fast-ViTs outperform 1D-CNNs, as each pixel contains only limited RGB bands, compared with the richer wavelength set found in HS data. In these cases, the lack of extensive spectral features needs to be compensated by incorporating neighboring spatial context.
V Conclusion
In this work, we have trained and tested 20 models, including 1D-CNNs, 2D-CNNs and vision transformers, for segmentation of hyperspectral satellite imagery with a focus on lightweight models for future in-orbit edge inference. We have proposed a 1D-CNN, called 1D-Justo-LiuNet, which has exceeded expectations by outperforming all other state-of-the-art models. It has outclassed conventional 2D-CNN UNets, their lighter versions, as well as recent vision transformers for edge inference. 1D-Justo-LiuNet has achieved top accuracy with the smallest model size among all tested models. Our results have also been validated across multiple satellite datasets. First, we used the hyperspectral mission HYPSO-1 as the primary case study for pixel-level segmentation. We then conducted further generalization tests on NASA’s EO-1 hyperspectral mission, and obtained consistent results. We additionally evaluated the model on RGB Satellite Imagery of Dubai, where we confirmed our findings. We conclude that 1D-CNNs, specifically 1D-Justo-LiuNet and 1D-Justo-HuNet, are highly suitable for future in-orbit deployment. As future work, we propose deploying 1D-Justo-LiuNet in flight for optimal satellite operations.
@@ V Conclusion
The labeled dataset from HYPSO-1 used in this paper is openly available to download at https://ntnu-smallsat-lab.github.io/hypso1_sea_land_clouds_dataset/. For any issues with downloading the dataset, contact NTNU’s HYPSO team. Additionally, our supplementary material is openly available at https://github.com/NTNU-SmallSat-Lab/s_l_c_segm_hyp_img/. Here, we provide the following resources. Firstly, the Python model codes for the 1D-CNNs, 2D-CNNs, and a reference to Apple’s FastViT. The codes are accompanied by a notebook that illustrates how to perform inference on HYPSO-1 data, by 1D-CNN and 2D-CNN models. Secondly, there is an executable software program designed to serve as an oceanic-terrestrial-cloud segmenter for hyperspectral imagery acquired by the HYPSO-1 satellite. Thirdly, we provide trained models with learned parameters, with 46 distinct configurations, for the HYPSO-1 data. Fourth, we offer additional in-depth architectural details for the models. Fifth, we include information regarding the training, validation, and test data splits. Sixth, we share an extensive set of experimental figures, including multiple metrics. Seventh, we provide a deployment set consisting of 30 unlabeled HYPSO-1 images. Lastly, we present the generalization tests in the complete deployment set.
Acknowledgement
We extend our gratitude to Anum Masood, Per Gunnar Kjeldsberg, and Simen Netteland.
@@ Appendix A Direct Channel Reduction
We obtain another HYPSO-1 dataset by reducing the 112 channels to merely 3 channels. To achieve this reduction, we explore the literature of dimensionality reduction extensively studied for HS satellite missions [53, 54, 55]. First, the focus on reducing dimensions originates from the need of having low-dimensional data to save computation resources, memory and power. Techniques such as Principal Component Analysis (PCA) [56] and Independent Component Analysis (ICA) [57] are proposed in the literature. In alignment with the state of the art, we employ PCA to reduce the 112 channels (normalized) to 3. To alleviate to a certain extent the computational demands of satellites operating with tight power constraints, we propose to conduct both the training and inference of PCA exclusively on ground. Eliminating PCA inference on the satellite enables potential power savings, which could be of importance to conduct the subsequent segmentation inference during flight. Therefore, and as also outlined by the literature [12], we suggest implementing a direct channel selection (3 channels) on board, instead of deploying PCA inference. In this work, the purpose of employing PCA only on ground is to identify in advance the channels containing the most useful information. In particular, we use PCA to identify the channels that have the greatest influence on a single principal component, as determined by the weights of those channels. For the 112-channel HYPSO-1 dataset in L1b radiance, we find that the majority of influential channels are predominantly within the NIR range. However, given the necessity of considering other spectral ranges in addition to NIR for the adequate segmentation of sea, land, and clouds – as the NIR range alone is likely to be insufficient – we opt to pick the top three influential channels from the visible blue, the combined visible green and red, and finally, the NIR. In the dataset, the most influential channels within each spectral range are: 412.72 nm (channel 7) for the blue spectrum, 699.61 nm (channel 89) in the green and red spectra, and 747.77 nm (channel 103) in the NIR spectrum. These closely align with the bands often analyzed during HYPSO-1 operations, especially for channels 89 and 103.